# iMAD: Intelligent Multi-Agent Debate for Efficient and Accurate LLM Inference

### v1

@@ 5.2 Main Results
We evaluate the token efficiency and accuracy of iMAD across six diverse datasets, comparing it against three categories of baselines: single-agent methods (CoT and SC), full-debate MAD frameworks (MAD and GD), and a confidence-based selective MAD approach (DOWN).
Advantage over Single-Agent and Full-Debate MAD Baselines.
As shown in Table 3, iMAD achieves superior token efficiency while maintaining or improving accuracy over all single-agent and full-debate MAD baselines. While CoT uses the fewest token, iMAD achieves up to 13.5% higher accuracy. Compared to SC, iMAD drastically reduces token usage with consistently higher accuracy. For example, on MEDQA, iMAD reduces token cost by 62.7% while improving accuracy by 4.7%. Against full-debate MAD baselines, iMAD reduces token usage significantly while achieving comparable or higher accuracy. On MEDQA, iMAD uses 68% fewer tokens than MAD and 92% fewer than GD, while achieving the highest accuracy among the three. A key to iMAD’s zero-shot efficiency is the classifier’s strong generalization: we trained it on two representative datasets selected to capture diverse model reasoning behaviors, enabling it to learn generalizable model behavior. The inefficiency of the full-debate strategy is especially evident on GSM8K, where iMAD outperforms MAD by 8.4% in accuracy. The only exception is MMLU, where iMAD slightly underperforms MAD. As shown in Table 4, we skip 3.5% of necessary debate, slightly higher than the 3.4% where a triggering debate corrects the error. Many questions in MMLU are short, factual, and span diverse domains, so incorrect single-agent answers often sound fluent and highly confident, with few hesitation cues. As a result, the debate classifier is more likely to skip debates that are actually needed.
@@
As shown in Table 3, DOWN and iMAD incur comparable token costs, with DOWN using slightly fewer tokens by omitting self-critique and skipping some needed debates. However, it consistently underperforms both iMAD and full-debate MAD in accuracy as it cannot tune thresholds in the zero-shot setting. In contrast, iMAD incurs a slight token overhead due to self-critique prompting and more necessary debates, but this cost is justified by improved identification of recoverable incorrect answers. For example, on OKVQA, DOWN’s accuracy remains near the single-agent baseline, revealing its inability to identify when debate is truly needed. This limitation arises because DOWN learns data-specific model behavior rather than generalizable behavior. In contrast, iMAD captures hesitation cues through a classifier trained on diverse features, enabling more accurate debate decisions and achieving higher accuracy with fewer tokens in zero-shot settings.
@@ 5.3 Breakdown of iMAD Debate Decisions
To evaluate the effectiveness of iMAD’s selective debate triggering, we analyze its decision outcomes across all datasets in Table 4. For each instance, we precompute both the single-agent and MAD outputs to establish whether MAD improves the answer, serving as the ground truth for evaluating decisions. We then measure how often iMAD is aligned with these beneficial outcomes. Overall, up to 95.9% of iMAD’s decisions are beneficial. When skipping debate, iMAD preserves correct answers ($\rightarrow$✓) in 65-80% of cases and avoids wasted computation on unrecoverable errors ($\rightarrow$✗) up to 13%. When triggering debate, iMAD often recover incorrect answers ($\rightarrow$✓). For example, iMAD successfully flips 16.2% of cases on GSM8K and 7.1% on MEDQA, approaching their respective upper bounds of 19.1% and 11.9% (see Table 2). Crucially, harmful decisions, such as overturning correct answers ($\rightarrow$✗) or incurring unnecessary debate overhead ($\rightarrow$✗ and $\rightarrow$✓), remain consistently low (typically around 5-10%). These results highlight iMAD’s ability to selectively trigger MAD only when it is likely to improve accuracy, while avoiding unnecessary token costs.
@@ 6 Conclusion
We presented iMAD, a token-efficient MAD framework that selectively triggers debates only when they are likely to improve outcomes. It combines structured self-critique prompting with a lightweight debate-decision classifier trained using our proposed FocusCal loss, enabling effective debate decisions based on single-agent responses without using evaluation data. Compared with five diverse baselines, iMAD reduces token usage by up to 92% while improving accuracy by up to 13.5%. These results show that iMAD is a practical and scalable solution for collaborative reasoning in agentic LLM systems. Future work includes exploring adaptive or online learning approaches to reduce labeling costs during classifier training and further improve generalization, as discussed in the extended version (see Appendix D).
@@ Acknowledgments
This research was supported in part by NSF grant CNS-2315851, the Commonwealth Cyber Initiative (CCI), and a Virginia Tech Presidential Postdoctoral Fellowship.
@@
We further include implementation details, additional analysis, additional results, and discussion in the appendix. First, we provide the exact prompt templates for all agents in QA and VQA tasks, including shared context, memory synchronization, role definitions and outputs, structured JSON formats, and placeholder fields in Appendix A. We also summarize all 41 interpretable features used in our debate decision classifier. In Appendix B, we present additional analysis on token efficiency and inference time, and we report ablation studies on the structured self-critique prompt and FocusCal loss, as well as cross-LLM results using Qwen 3.0 and GPT-5 nano in Appendix C. Lastly, we discuss our assumptions, limitations, and future work in Appendix D.
@@ Appendix A Implementation Details
A.1 Prompt Templates for VQA and QA Tasks
This section specifies the full prompt protocol used in our framework for VQA and QA tasks. We first describe how MAD maintains shared context and memory synchronization across agents, the roles and outputs of each agent, the required structured JSON format used by the Judge, and the placeholders that instantiate prompts at runtime. We then list the exact task-specific prompt templates for VQA and QA, specifying for each agent what it receives (inputs), the instruction it follows, and the required output.
@@ A.1 Prompt Templates for VQA and QA Tasks
When MAD is triggered, all agents read the same dialogue history. At the beginning of each round, we share the running transcript with every agent, which includes the question, the initial reasoning or self critique, and all prior messages from all agents, so that decisions are made with a shared context. This synchronization keeps the discussion coherent.
Roles and Outputs.
We use three roles in MAD: Debater (Affirmative), Debater (Negative), and Judge, plus a single agent baseline and a self critique stage in iMAD. Debaters must select exactly one option and provide a concise justification grounded in the available evidence (image for VQA, text for QA). The Judge returns structured JSON so that its decision can be parsed automatically, while the self-critique single agent produce free form text that is later featurized for the debate decision classifier.
@@
At the end of each round, the Judge returns a single JSON object with the fields {"Preference", "Supported Side", "Reason", "debate_answer"}. The system uses this object to control the debate: if "Preference" is "No", we start the next round; if it is "Yes", we stop the debate and record "debate_answer" when provided. For finalization, the Judge returns a JSON output with {"Reason", "debate_answer"}. We reject malformed JSON outputs such as extra text, missing keys, or invalid values and then reissue the prompt.
@@
We fill each template at runtime with concrete values drawn from the example and the current debate state. Example fields: {QUESTION}, {OPTIONS} (current option set), {OPTIONS2} (optional reduced set for the final decision), and {IMAGE} for VQA. Debate state fields: {ROUND} (current round index), {TRANSCRIPT} (the synchronized dialogue history), {AFF_ANS} and {NEG_ANS} (the latest messages from each side), and {OPP_ANS} (the opponent’s most recent message). Before each agent is prompted, we compute these values from the input example and the live transcript and then substitute them into the template.
@@
We present the QA prompt templates in execution order. The self critique single agent receives the question and the option set, produces an answer with reasons, and provides a concise self critique. When MAD is triggered, the Debater (Affirmative) selects one option and justifies it, and the Debater (Negative) challenges the affirmative position by arguing for a different option with a brief rationale. Both debaters read the synchronized transcript so decisions are made on a shared context. Finally, the Judge returns structured JSON that either requests another round or finalizes the answer by providing debate_answer and a short reason.
@@
Self-Critique Single-Agent Baseline Receives: {QUESTION}, {OPTIONS}. Instruction: Answer the multiple-choice question. Use a step-by-step chain of thought to justify your initial answer. Then, argue against your own answer with a plausible counter-reasoning — this is not optional. Finally, confirm your final answer and report confidence scores for both the initial belief and the forced disagreement. Output: One option (A/B/C/…); a short justification.
@@
Debater (Affirmative) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; initial stance {AFF_ANS}; and {OPP_ANS} from round 2+. Instruction: First turn — restate {AFF_ANS} and justify. Later turns — agree or refute {OPP_ANS}; if refuting, choose an alternative strictly from {OPTIONS} and justify. Output: A short argument plus one choice from {OPTIONS}.
@@
Debater (Negative) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; latest affirmative message {AFF_ANS}; {OPP_ANS}. Instruction: Disagree when warranted. Provide an alternative strictly from {OPTIONS} and justify succinctly. Output: A counter-argument plus one choice from {OPTIONS}.
@@
Judge (per round) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; {AFF_ANS}, {NEG_ANS}. Instruction: Evaluate accuracy and justification quality; determine preference and tentative answer if any. Output (JSON only): {”Preference”:”Yes—No”,”Supported Side”:”Affirmative—Negative”,”Reason”:”…”,”debate_answer”: one of {OPTIONS} or ””}.
@@
Judge (Finalization) Receives: {TRANSCRIPT}; {QUESTION}, {OPTIONS}; candidate list; {OPTIONS2}. Instruction: (A) List candidates strictly from {OPTIONS}. (B) Provide the final answer with a brief reason. Output (JSON): {”Reason”:”…”,”debate_answer”: one of {OPTIONS2}}.
@@
We present the VQA prompt templates in execution order. The self critique single agent receives the image, the question, and the option set, produces an answer with reasons grounded in visual evidence, and provides a concise self critique. When MAD is triggered, the Debater (Affirmative) selects one option and justifies it using concrete visual cues such as objects, attributes, or regions in the image, and the Debater (Negative) challenges the affirmative position by arguing for a different option with a brief image-grounded rationale. Similar to the QA task, both debaters read the synchronized transcript so decisions are made on a shared context. The Judge follows the same protocol: it returns structured JSON that either requests another round or finalizes the answer by providing debate_answer and a brief reason, with an explicit check for consistency with the image.
@@
Self-Critique Single-Agent Baseline Receives: {IMAGE}, {QUESTION}, {OPTIONS}. Instruction: From the image, answer the multiple-choice question. Use a step-by-step chain of thought to justify your initial answer. Then, argue against your own answer with a plausible counter-reasoning — this is not optional. Finally, confirm your final answer and report confidence scores for both the initial belief and the forced disagreement. Output: One option (A/B/C/…); a short justification.
@@
Debater (Affirmative) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; initial stance {AFF_ANS}; and {OPP_ANS} from round 2+. Instruction: First turn — restate {AFF_ANS} and justify from the image. Later turns — agree or refute {OPP_ANS}; if refuting, select an alternative strictly from {OPTIONS} and justify briefly. Output: A short argument plus one choice from {OPTIONS}.
@@
Debater (Negative) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; latest affirmative message {AFF_ANS}; {OPP_ANS}. Instruction: Disagree when warranted. Provide an alternative strictly from {OPTIONS} and justify using visual evidence or logic. Output: A counter-argument plus one choice from {OPTIONS}.
@@
Judge (per round) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; {AFF_ANS}, {NEG_ANS}. Instruction: Evaluate (1) accuracy w.r.t. image and question and (2) justification quality. Decide if there is a clear preference and provide a tentative answer if any. Output (JSON only): {”Preference”:”Yes—No”,”Supported Side”:”Affirmative—Negative”,”Reason”:”…”,”debate_answer”: one of {OPTIONS} or ””}.
@@
Judge (Finalization) Receives: {TRANSCRIPT}; {IMAGE}, {QUESTION}, {OPTIONS}; possibly reduced {OPTIONS2}. Instruction: (A) List candidate answers strictly from {OPTIONS}. (B) Provide the final decision with a brief reason. Output (JSON): {”Reason”:”…”,”debate_answer”: one of {OPTIONS2}}.
Table 5: List of extracted 41 features for the debate decision classifier.
@@
Question Surface-level statistics QuestionToken Number of tokens in the question
AnswerToken Number of tokens in the model’s final answer
question_named_entity_count Number of named entities (e.g., person, location, date)
@@
Part-of-speech counts Question_num_nouns Number of nouns in the question
Question_num_verbs Number of verbs in the question
Question_num_adjs Number of adjectives in the question
@@
InitialReason_HedgeCount Number of hedge words (e.g., “maybe”, “likely”)
InitialReason_CertaintyCount Number of certainty words (e.g., “definitely”, “certainly”)
InitialReason_contrast Number of contrastive discourse markers (e.g., “but”, “however”)
@@
Part-of-speech counts InitialReason_num_nouns Number of nouns in the justification
InitialReason_num_verbs Number of verbs in the justification
InitialReason_num_adjs Number of adjectives in the justification
@@
NegativeReason_HedgeCount Number of hedge words in the critique
NegativeReason_CertaintyCount Number of certainty words in the critique
NegativeReason_contrast Number of contrastive discourse markers in the critique
@@
Part-of-speech counts NegativeReason_num_nouns Number of nouns in the critique
NegativeReason_num_verbs Number of verbs in the critique
NegativeReason_num_adjs Number of adjectives in the critique
A.2 Extracted Features for Debate Decision Classification
In this section, we provide a detailed description of the features that drive the debate decision classifier and evaluate their importance and necessity.

### v2

@@ 5.1 Main Results
We compare the token efficiency and accuracy of iMAD against single-agent, full-debate MAD, and selective MAD.
Advantage over Single-Agent and Full-Debate MAD.
As shown in Table 3, iMAD achieves superior token efficiency while maintaining or improving accuracy over all single-agent and full-debate MAD baselines. While CoT uses the fewest tokens, iMAD achieves up to 13.5% higher accuracy. Compared to SC, iMAD drastically reduces token usage with consistently higher accuracy. For example, on MEDQA, iMAD reduces token cost by 62.7% while improving accuracy by 4.7%. Against full-debate MAD baselines, iMAD reduces token usage significantly while achieving comparable or higher accuracy. On MEDQA, iMAD uses 68% fewer tokens than MAD and 92% fewer than GD, while achieving the highest accuracy. A key to iMAD’s zero-shot efficiency is the classifier’s strong generalization: we trained it on two representative datasets to capture diverse model reasoning behaviors, enabling it to learn generalizable model behaviors. iMAD’s advantage over the full-debate strategy is especially evident on GSM8K, where iMAD outperforms MAD by 8.4% in accuracy. The only exception is MMLU, where MAD performs slightly better. In Table 4, the classifier skips debate in 3.5% of questions where it would help and 3.4% where a triggered debate fixes an error. This is because many MMLU questions are short and factual across diverse domains, wrong single-agent answers often sound fluent and confident, giving few hesitation cues and causing the classifier to miss some needed debates.
@@
As shown in Table 3, DOWN and iMAD incur comparable token costs, with DOWN using slightly fewer tokens by omitting self-critique and skipping some needed debates. However, both iMAD and MAD consistently achieve higher accuracy than DOWN, since DOWN cannot tune its confidence thresholds in the zero-shot setting. In contrast, iMAD incurs slightly more tokens due to self-critique prompting and more necessary debates, but this cost is justified by improved identification of recoverable errors. For example, on OKVQA, DOWN’s accuracy remains near the single-agent baseline, revealing its inability to identify when debate is truly needed. This limitation arises because DOWN learns data-specific model behavior rather than generalizable behavior. In contrast, iMAD captures hesitation cues through a classifier trained on diverse features, enabling more accurate debate decisions and achieving higher accuracy with fewer tokens in zero-shot settings.
@@ 5.2 Breakdown of iMAD Debate Decisions
To evaluate the effectiveness of iMAD’s selective debate triggering, we analyze its decision outcomes across all datasets in Table 4. For each instance, we precompute the single-agent and MAD outputs to establish whether MAD improves the answer, serving as the ground truth for evaluating decisions. We then measure how often iMAD matches the beneficial outcomes. Overall, up to 95.9% of iMAD’s decisions are beneficial. When skipping debate, iMAD preserves correct answers ($\rightarrow$✓) in 65-80% of cases and avoids wasted computation on unrecoverable errors ($\rightarrow$✗) by up to 13%. When triggering debate, iMAD often recover incorrect answers ($\rightarrow$✓). For example, iMAD successfully flips 16.2% of cases on GSM8K and 7.1% on MEDQA, approaching their respective upper bounds of 19.1% and 11.9% (see Table 2). Crucially, harmful decisions, such as overturning correct answers ($\rightarrow$✗) or incurring unnecessary debate overhead ($\rightarrow$✗ and $\rightarrow$✓), remain consistently low (around 5-10%). These results highlight iMAD’s ability to selectively trigger MAD only when it is likely to improve accuracy, while avoiding unnecessary token costs.
@@ 6 Conclusion
We presented iMAD, a token-efficient MAD framework that triggers debates only when it is likely to improve outcomes. It combines structured self-critique prompting with a lightweight debate decision classifier trained with FocusCal loss, enabling effective debate decisions based on single-agent responses without using evaluation data. Compared with five baselines, iMAD reduces token usage by up to 92% while improving accuracy by up to 13.5%. These results show that iMAD is a practical and scalable solution for collaborative reasoning in agentic LLM systems. Future work includes exploring adaptive or online learning approaches to reduce labeling costs during classifier training and further improve generalization, as discussed in Appendix D.
@@ Acknowledgments
We sincerely thank the anonymous reviewers of the AAAI’26 for their constructive comments and suggestions. This research was supported in part by NSF grant CNS-2315851, the Commonwealth Cyber Initiative (CCI), and Virginia Tech Presidential Postdoctoral Fellowship.
@@
In the appendix, we include implementation details, additional performance analyses, additional evaluation results, and discussions. First, we provide the prompt templates and 41 features in our debate decision classifier in Appendix A. Then, we present additional performance analyses on token efficiency and inference time in Appendix B and report ablation studies on the structured self-critique prompt and FocusCal loss, as well as cross-LLM results using Qwen 3.0 and GPT-5 nano in Appendix C. Lastly, we discuss our assumptions, limitations, and future work in Appendix D.
@@ Appendix A Implementation Details
We provide the exact prompt templates for all agents in QA and VQA tasks, including shared context, memory synchronization, role definitions and outputs, structured JSON formats, and placeholder fields in Appendix A.1, and summarize all 41 interpretable features used in our debate decision classifier in Appendix A.2.
A.1 Prompt Templates
This section specifies the full prompt protocol used in our framework. We first introduce the agent roles in MAD, including how the Debater and Judge agents interact during the debate and what outputs are expected from each agent. We then describe how MAD maintains shared context and memory synchronization across these agents. In contrast, Debater agents are not required to follow any structured format and instead produce free-form natural language responses. We then introduce the prompt placeholders that are filled in at runtime with instance-specific values and the current debate state, so that each agent receives a prompt that reflects the correct question, option set, and transcript. Finally, we list the exact task-specific prompt templates for VQA and QA, that specify what input each agent receives, the instruction it follows, and the output it generates.
Roles and Outputs.
Prior to deciding whether to trigger MAD, iMAD first executes a self-critique stage with a single self-critique agent. This self-critique single agent produces a preliminary answer together with a step-by-step explanation and a mandatory self-critique. The debate decision classifier then takes this self-critique response as input and decides whether a debate is necessary. If MAD is triggered, the framework instantiates three distinct agent roles: the Debater (Affirmative) agent, the Debater (Negative) agent, and the Judge agent. The Debater (Affirmative) agent takes the single-agent self-critique answer as its initial position and provides reasons and evidence to support this answer. In contrast, the Debater (Negative) agent reviews the affirmative reasoning and highlights potential issues. When the Debater (Negative) agent disagrees, it proposes an alternative option and provides its own justification. During each debate round, both debaters are required to select exactly one option from the candidate set and provide a concise rationale grounded in the given context (e.g., the visual content for VQA or the textual passage for QA). The Judge agent then assesses the arguments from both sides and produces a structured JSON output that tells the system whether to proceed to another debate round or to terminate the process and return the final answer.
@@ A.1 Prompt Templates
When MAD is triggered, at the beginning of each debate round, the system constructs a running transcript that includes the question, the initial CoT reasoning and self-critique, and all previous messages from every agent. This transcript is then inserted into the prompt of each agent, so they all see the same question, reasoning, and prior messages before generating their next response. In other words, before each debate round, we synchronize the memory of all agents by giving them the same transcript. This memory synchronization ensures that agents always make decisions under a consistent context and keeps the multi-round discussion coherent.
@@
At the end of each debate round, the Judge agent generates a single, parsable JSON output containing four specific fields: Preference, Supported Side, Reason, and Debate Answer. The Preference field takes values “Yes” or “No” and controls whether another debate round is started. If Preference is set to “No” and the current number of rounds is still below a preset maximum number of debate rounds, the Judge is indicating that neither the affirmative nor the negative side has presented a sufficiently convincing argument to settle the dispute. Consequently, the system initiates the next round of debate to solicit further reasoning. We manually set this maximum number of debate rounds to a small constant to keep the token cost under control. In all our experiments, we set the debate at 5 rounds. If the maximum number of rounds is reached without the Judge ever setting Preference to “Yes”, the system enters a finalization stage in which the Judge must select the best-supported answer from the existing arguments and produce a final decision.
Conversely, if Preference is set to “Yes”, the Judge has identified a clear superior argument from a preferred side, and the Supported Side field is then set to either “Affirmative” or “Negative”. In this case, the debate terminates immediately. The system records the content of the Debate Answer field as the judge’s current best answer selected from the option set, and the Reason field provides a brief natural language explanation of why this side and answer are preferred. For the finalization stage (e.g., after the maximum number of rounds is reached), the Judge produces a simplified JSON output containing only the Reason and the final Debate Answer. To make the control logic robust, we validate every JSON output returned by the Judge agent. We mark a JSON output as malformed whenever it has text outside the JSON object or misses required keys. We also reject outputs that use invalid field values (e.g., a Preference other than “Yes” or “No”) or specify a Debate Answer that is not in the option set. In such cases, we discard the response and reissue the Judge prompt until a valid JSON output is obtained.
@@
In our prompt templates, we use placeholders as dynamic fields that are filled in at runtime to adapt the instructions to a specific question and debate round. Each placeholder is a symbolic token, such as {QUESTION} or {ROUND}, which will be replaced by the system with a concrete value before sending the prompt to an agent. For example, {QUESTION} is replaced by the actual question text, {OPTIONS} by the current multiple-choice option set, and {TRANSCRIPT} by the full history of messages in the current debate. All agents use these placeholders in their prompt templates, so each prompt reflects the current question, its multiple-choice option set, the up-to-date debate transcript, and other instance-specific information.
We group these placeholders into two categories: example fields and debate state fields. The example fields describe the static information for a given instance. We use {QUESTION} for the question text, {OPTIONS} for the current multiple-choice option set of that question, {OPTIONS2} for a reduced option set constructed from {OPTIONS} by collecting all distinct options proposed by any agent during the debate and using this subset only in the final decision step, and {IMAGE} to refer to the visual input in VQA tasks. These fields do not change across debate rounds for the same example. The debate state fields describe the dynamic status of the ongoing debate. We use {ROUND} to record the current round index and {TRANSCRIPT} to store the synchronized dialogue history, including the question, the initial reasoning or self-critique, and all previous messages from every agent. We further use {AFF_ANS} and {NEG_ANS} to denote the latest messages from the affirmative and negative debaters, respectively. From the perspective of the current speaker, {NEG_ANS} refers to the opponent’s most recent message.
Before each agent is prompted, the system first computes the current values of all relevant placeholders from the input example and the up-to-date transcript. The system then substitutes these values into the template. This placeholder substitution ensures that every agent receives a prompt that accurately reflects both the fixed information about the instance and the current state of the debate.
@@
We use a different prompt template for each agent and execute them in a fixed order. We first apply the Self-Critique Single Agent template. In this template, the agent receives {QUESTION} and {OPTIONS}, and is asked to choose an initial answer, explain it step by step, then deliberately argue against this answer, and finally state a final choice together with confidence scores for both the initial belief and the self-critique. The entire response is passed to the debate decision classifier. If the classifier decides to trigger MAD, we then use the two debater templates. The Debater (Affirmative) template takes the synchronized transcript {TRANSCRIPT}, the question and option set, and the current affirmative stance {AFF_ANS}. In the first debate round, it restates {AFF_ANS} and provides supporting reasons. In later rounds, it also reads {NEG_ANS} and either maintains its current answer or switches to a different option from {OPTIONS}, with a brief justification in each case. The Debater (Negative) template also reads {TRANSCRIPT}, the question, and the option set. It examines the affirmative message {AFF_ANS}, highlights potential issues, and when it disagrees, selects another option from {OPTIONS} and provides its own justification. Finally, we use two Judge templates. The per-round Judge template asks the Judge agent to compare {AFF_ANS} and {NEG_ANS}, decide whether one side is clearly better, and return a structured JSON output, which the system uses to either start another round or stop the debate. The finalization Judge template is then used to list plausible candidates and commit to a single Debate Answer from {OPTIONS2} with a short reason to explain the decision.
@@
Self-Critique Single Agent Input: {QUESTION} and {OPTIONS}. Instruction: Answer the multiple-choice question. Use a step-by-step chain of thought to justify your initial answer. Then, argue against your own answer with a plausible counter-reasoning — this is not optional. Finally, confirm your final answer and report confidence scores for both the initial belief and the forced disagreement. Output: One option from (A/B/C/…), the initial reasoning with confidence score, and the self-critique reasoning with confidence score.
@@
Debater (Affirmative) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {QUESTION}, {OPTIONS}, and {AFF_ANS} (affirmative agent’s fixed initial stance, which is passed in every round to remind the agent of its side). Instruction: First turn — restate {AFF_ANS} and justify. Later turns — agree or refuse {NEG_ANS}; if refusing, choose an alternative strictly from {OPTIONS} and justify. Output: A short argument and one selected option from {OPTIONS}.
@@
Debater (Negative) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {QUESTION}, {OPTIONS}, and latest affirmative message {AFF_ANS} for the current round. Instruction: Disagree when warranted. Provide an alternative strictly from {OPTIONS} and justify succinctly. Output: A short counter-argument (or explicit agreement) and one selected option from {OPTIONS}.
@@
Judge (Per Round) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {QUESTION}, {OPTIONS}, and {AFF_ANS} and {NEG_ANS} for the current round. Instruction: Evaluate accuracy and justification quality; determine preference and tentative answer, if any. Output (JSON): {Preference: “Yes||No”, Supported Side: “Affirmative||Negative”, Reason: …, Debate Answer: one of {OPTIONS}}.
@@
Judge (Finalization) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {QUESTION}, {OPTIONS}, and reduced candidate list {OPTIONS2}. Instruction: (A) List candidates strictly from {OPTIONS}. (B) Provide the final answer with a brief reason. Output (JSON): {Reason: …, Debate Answer: one of {OPTIONS2}}.
@@
For VQA tasks, we reuse the same agent roles and control flow as in QA tasks, but extend every template to condition on the image and to explicitly mention visual evidence in the instructions. The Self-Critique Single Agent template now takes {IMAGE}, {QUESTION}, and {OPTIONS} as input and begins with the instruction “From the image, answer the multiple-choice question,” so that both the initial answer and the self-critique are written with the image in view. The Debater (Affirmative) and Debater (Negative) templates also receive {IMAGE} together with the synchronized transcript, and their instructions explicitly ask the debaters to “justify from the image” or to “justify using visual evidence or logic” when agreeing, disagreeing, or proposing a new option. In practice, this encourages debaters to support their positions by referring to concrete visual cues in the image (e.g., objects, attributes, or spatial relations). Finally, the per-round and finalization Judge templates keep the same JSON structure as in QA, but now take {IMAGE} as an additional input and explicitly instruct the Judge to “evaluate accuracy w.r.t. image and question” and to ensure that the preferred side and final Debate Answer are consistent with both the image and the debate dialogue history.
@@
Self-Critique Single Agent Input: {IMAGE}, {QUESTION}, and {OPTIONS}. Instruction: From the image, answer the multiple-choice question. Use a step-by-step chain of thought to justify your initial answer. Then, argue against your own answer with a plausible counter-reasoning — this is not optional. Finally, confirm your final answer and report confidence scores for both the initial belief and the forced disagreement. Output: One option (A/B/C/…), the initial reasoning with confidence score, and the self-critique reasoning with confidence score.
@@
Debater (Affirmative) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {IMAGE}, {QUESTION}, {OPTIONS}, and {AFF_ANS} (affirmative agent’s fixed initial stance, which is passed in every round to remind the agent of its side). Instruction: First turn — restate {AFF_ANS} and justify from the image. Later turns — agree or refuse {NEG_ANS}; if refusing, select an alternative strictly from {OPTIONS} and justify briefly. Output: A short argument and one selected option from {OPTIONS}.
@@
Debater (Negative) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {IMAGE}, {QUESTION}, {OPTIONS}, and latest affirmative message {AFF_ANS} for the current round. Instruction: Disagree when warranted. Provide an alternative strictly from {OPTIONS} and justify using visual evidence or logic. Output: A short counter-argument (or explicit agreement) and one selected option from {OPTIONS}.
@@
Judge (Per Round) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {IMAGE}, {QUESTION}, {OPTIONS}, and {AFF_ANS} and {NEG_ANS} for the current round. Instruction: Evaluate (1) accuracy w.r.t. image and question and (2) justification quality. Decide if there is a clear preference and provide a tentative answer if any. Output (JSON): {Preference: “Yes||No”, Supported Side: “Affirmative||Negative”, Reason: …, Debate Answer: one of {OPTIONS}}.
@@
Judge (Finalization) Input: {TRANSCRIPT} (running log includes the initial single-agent answer, {AFF_ANS} and {NEG_ANS} choices with their arguments, and the Judge’s outputs from all previous debate rounds), {IMAGE}, {QUESTION}, {OPTIONS}, and reduced candidate list {OPTIONS2}. Instruction: (A) List candidate answers strictly from {OPTIONS}. (B) Provide the final decision with a brief reason. Output (JSON): {Reason: …, Debate Answer: one of {OPTIONS2}}.
@@
Input Question Surface-level statistics QuestionToken # of tokens in the question
AnswerToken # of tokens in the model’s final answer
question_named_entity_count # of named entities (e.g., person, location, date)
@@
Part-of-speech counts Question_num_nouns # of nouns in the question
Question_num_verbs # of verbs in the question
Question_num_adjs # of adjectives in the question
@@
InitialReason_HedgeCount # of hedge words (e.g., “maybe”, “likely”)
InitialReason_CertaintyCount # of certainty words (e.g., “definitely”, “certainly”)
InitialReason_contrast # of contrastive discourse markers (e.g., “however”)
@@
Part-of-speech counts InitialReason_num_nouns # of nouns in the justification
InitialReason_num_verbs # of verbs in the justification
InitialReason_num_adjs # of adjectives in the justification
@@
NegativeReason_HedgeCount # of hedge words in the critique
NegativeReason_CertaintyCount # of certainty words in the critique
NegativeReason_contrast # of contrastive discourse markers in the critique
@@
Part-of-speech counts NegativeReason_num_nouns # of nouns in the critique
NegativeReason_num_verbs # of verbs in the critique
NegativeReason_num_adjs # of adjectives in the critique
Table 5: List of extracted 41 features used in the debate decision classifier.
A.2 Features for Debate Decision Classification
In this section, we provide a detailed description of the features used by the debate decision classifier and evaluate their importance using SHAP and PCA. We then assess the necessity of features by removing the bottom 20% of features based on a joint importance score that combines SHAP and PCA and comparing the accuracy and token costs with those of the full feature set.